# Enhancing Negation Awareness in Universal Text Embeddings: A Data-efficient and Computational-efficient Approach

### v1

@@ IV Improve negation awareness in universal text embeddings
From the analysis in the previous section, it can be concluded that the negation awareness of the state-of-the-art universal text embeddings is limited. Previous studies [41, 42, 43] assume that the embedding models lack of negation knowledge and focus on improving negation awareness by fine-tuning pre-trained models on negation datasets. However, fine-tuning on negation datasets may hurt the performance of universal text embeddings on other tasks (e.g. retrieval or topic modeling) which requires a better understanding of topic information. To efficiently deal with such conflicts, a data-efficient and sustainable embedding re-weighting function is proposed in this section without modifying the parameters of text embedding models. The main hypothesis is that universal text embeddings inherently encapsulate a broad but imbalanced spectrum of semantic information including topics, sentiment, negation and so on due to their large scale training/fine-tuning on extensive and diverse datasets. By assigning greater weights to the dimensions that predominantly capture negation-related information, they are expected to recognize and process negations more effectively. To ensure the universality across different tasks: this negation-aware weight will only be applied for negation related tasks, while the original unweighted embedding is used for other tasks.
@@
where $a$ is a hyperparameter to control the trade off between negation information and general semantic information in text embeddings. One toy example of the the average contribution of each dimension to negation awareness $\overline{V}$ is shown in Figure 3 (a), the corresponding dimension weights when applying Equation 4 are shown in Figure 3 (b), (c), (d) when $a$ equals to 2, 4, 6 respectively. It can be seen that larger $a$ assigns larger weights on to the dimension with positive contribution to negation awareness.
For negation related tasks, the dimension weights are then applied through element-wise multiplication to the original embeddings. This is equivalent to apply a weighted cosine similarity on the original embeddings.
Figure 3: A toy example of $\overline{V}$ with 100 dimensions is shown in (a): X-axis is the embedding dimension, Y-axis is the corresponding contribution to negation awareness. The impact of hyperparameter $a$ on the dimension weights are shown in (b), (c), (d) when $a$ equals to 2, 4, 6 correspondingly: X-axis is the dimension, Y-axis is the corresponding weight.
V Experimental results
V-A Experiments on simple negation
V-A1 Datasets
@@ V-A2 Experiment Setups
V-A3 Experimental results
TABLE I: Performances of text embeddings on STSB dataset: corr is the correlation between original embedding similarity and the human annotations, $corr\_W_{stsb}$ is the correlation between weighted embedding similarity and the human annotations; acc is the accuracy of original embedding; $acc\_W_{stsb}$ is the accuracy of weighted embedding.
@@ V-A3 Experimental results
The experimental results of state-of-the-art text embeddings on STSB dataset are shown in Table I: acc is the test accuracy of original embedding; $acc\_W_{stsb}$ is the accuracy of weighted embedding where the weight is calculated following Equation 4. Generally speaking, universal text embeddings have better performance than the two baseline contextual text embedding models. The proposed dimension weighting method improves the accuracy across all text embeddings with an average improvement of 4.51%. bge-large-en-v1.5 benefits the most from the proposed method with an improvement of 6.91% while gte-base-en-v1.5 has the smallest improvement of 0.91%. To study if improved negation awareness would hurt the topic information understanding, the correlation between human annotated similarity values (mainly based on topic information) and embedding semantic similarity (before and after applying weights) is also calculated as shown in Table I: corr is the correlation between original embedding similarity and the human annotations, $corr\_W_{stsb}$ is the correlation between weighted embedding similarity and the human annotations. On average, the semantic similarity based on weighted embeddings have slightly reduced correlation (1% less) with human annotations compared to the original embedding.
@@
V-B Generalization ability of weights learnt from STSB to SemAntoNeg (more complex negation)
@@
TABLE II: The the generalization ability of the embedding weights learnt from STSB on SemAntoNeg Benchmark.
@@
Compared to STSB data with simple negation, SemAntoNeg Benchmark is more difficult as it contains negation, antonym and double negation (negation plus antonym). In this section, we evaluate the generalization capability of the simple negation awareness embedding weights, learned from STSB, on the SemAntoNeg Benchmark. The results are presented in Table II. The overall performance of universal text embeddings (with the mean performance of 58.64%) are much better than that of contextual text embedding baselines (with the mean performance of 33.32%). The random guess performance on SemAntoNeg Benchmark is 33.33%, which indicates that contextual text embeddings is prone to predicting the sentence with the highest lexical overlap while ignoring the negation [38]. On average (across all text embedding models), the embedding weights learnt from STSB improves the performance of the original embedding by 2.41%. This suggests that the weights learnt from STSB are able to generalize to SemAntoNeg Benchmark. However, the performance gains for e5-large-v2 and gte-base-en-v1.5 are less noticeable. In terms of model parameter size, it is observed that larger universal text embeddings demonstrate superior comprehension of complex negation compared to their smaller counterparts.
V-C Generalization ability of proposed solution on SemAntoNeg Benchmark
TABLE III: The generalization ability of proposed solution on SemAntoNeg Benchmark: $acc\_W_{anto}^{K}$ means the performance of embedding weights learnt from K training samples (K = 200, 500, 100). The mean and standard deviations of accuracy evaluated over 10 runs are shown in the table.
@@ V-C Generalization ability of proposed solution on SemAntoNeg Benchmark
In this section, the objective is to study whether the proposed embedding re-weighting solution is able to learn from complex negation (negation + antonym) from SemAntoNeg Benchmark. Unlike STSB, SemAntoNeg Benchmark has no predefined train test split. Hence, a stratified repeated random sampling approach is used to achieve a robust estimate of the performance. The random splitting is repeated 10 times, with 32% as training data (1000x4 samples) and 68% as test data (2152x4 samples). To test the data efficiency of the proposed solution, we test 3 different training data sizes: 200, 500 and 1000. The corresponding test performances (mean and standard deviations of accuracy evaluated over 10 runs) are shown in Table III as $acc\_W_{anto}^{200}$, $acc\_W_{anto}^{500}$, and $acc\_W_{anto}^{1000}$ respectively.
Generally speaking, the proposed weighting method is able to improve the performance on complex negation tasks across all text embedding models with small size training data. When the training data size is only 200 ($acc\_W_{anto}^{200}$ in Table III), the universal text embeddings’ performance are improved by 14.52% on average, while the contextual text embeddings’ performance are improved by 11.29%. When increasing the training data size from 200 to 500, the average universal text embeddings’ performance are further improved slightly by 1.34% with reduced standard deviations, while the average contextual text embeddings’ performance are further improved by 0.97%. Wilcoxon-Holm post hoc test with Critical Differences (CD) is also done to have an overall statistical comparison as shown in Figure 4: the proposed solution with three training sizes are all significantly better than the original embeddings, but there is no significant difference between the performance on 500 training datasize and 1000 training datasize (when alpha is 0.00001). Similar to the observations in previous sections: 1. larger universal text embeddings demonstrate superior comprehension of negation compared to their smaller counterparts, 2. universal embeddings benefit more from the proposed solution than contextual embeddings.
@@
V-D Generalization to Large Language Model based universal text embeddings
@@
The objective in this section is to study whether LLM-based high dimensional task-specific embeddings can be further improved by the proposed dimension weighting method. Three state-of-the-art LLM-based text embeddings are selected: SFR-Embedding-2_R with Mistral-7B [47] as backbone, gte-Qwen2-7B-instruct with Qwen2-7B [48] as backbone and gte-Qwen2-1.5B-instruct with Qwen2-1.5B [48] as backbone. The experimental results are shown in Table IV: Similar to the observations from previous sections, the proposed data efficient solution can improve significantly (around 12% improvement on average) the negation awareness of LLM-based text embeddings with small training data size.
@@
TABLE IV: The performance of proposed solution with LLM-based embedding models on SemAntoNeg Benchmark: $acc\_W_{anto}^{K}$ means the performance of embedding weights learnt from K training samples (K = 200, 500, 100). The mean and standard deviations of accuracy evaluated over 10 runs are shown in the table.
@@ V-E Summary
In this section, STSB dataset with simple negation is used firstly to test the effectiveness of the proposed dimension re-weighting method. The proposed solution is able to improve the accuracy of negation awareness task of both BERT-based universal text embeddings (4.68% improvement) and contextual text embeddings (3.83% improvement). The weights learnt on STSB are subsequently applied to SemAntoNeg Benchmark with complex negation (negation + antonym). The experimental results show that the STSB learnt weights can also improve the complex negation awareness with over 2% improvement on BERT-based universal text embedding models. Furthermore, the proposed solution is used to learn weights from small training data on SemAntoNeg Benchmark (200, 500 and 1000): performances based on all three training datasizes are significantly better than the original embedding (with over 10% improvement) across all BERT-based embedding models, which shows that the proposed solution is robust and data efficient. Finally, we assess the method’s effectiveness on three LLM-based embeddings. The results demonstrate that the proposed method can also improve the negation-awareness of high dimensional task-specific LLM-based universal embeddings.
Sustainability: The environmental impact of AI solutions has attracted a lot of attention in recent years [49, 50]. Compared to existing solutions in the literature, the proposed solution is significantly more sustainable in terms of training datasize, training time and computational cost.
VI Conclusion
@@
Recent literature has witnessed the fast progress and development of universal text embeddings. However, due to the bias in popular evaluation benchmarks, the negation awareness capacity of these models remains unclear. In this work, we start with a holistic analysis in order to answer how similar are negated text pairs interpreted by the state-of-the-art universal text embeddings. The experimental results show that the state-of-the-art universal text embeddings lack negation awareness and interpret negated text pairs as roughly similar in terms of semantic meaning. To efficiently deal with the conflict that different tasks need different trade-off between topic and negation information among other semantic information, a data-efficient and sustainable embedding re-weighting function is proposed without modifying the parameters of text embedding models. The proposed solution is able to improve text embedding models’ negation awareness significantly on both STSB with simple negations and SemAntoNeg with complex negations. Furthermore, the proposed solution can also improve the negation awareness of LLM-based task-specific high dimensional embeddings. An additional benefit of the proposed solution is its flexibility: re-weighting is applied exclusively to tasks that need strong negation awareness, while other tasks like topic clustering/classification use the original unweighted embedding, which ensures the universality of universal text embeddings across different tasks. Future work will focus on generalizing the proposed solution to other NLP tasks requiring trade-offs among different types of semantic information. Another research direction is to combine the proposed data efficient solution with automatic training data generation by LLMs in order to improve the generalization ability of universal text embedding models on unseen tasks.

### v2

@@ 4 Negation Adapter: A Lightweight Alternative to Fine-Tuning
To deal with these limitations, we propose a lightweight, post-hoc re-weighting “adapter" which leaves the base model intact and only adjusts its use when negation sensitivity is required. The main hypothesis is that universal text embeddings inherently encapsulate a broad but imbalanced spectrum of semantic information including topics, sentiment, negation and so on due to their large scale training/fine-tuning on extensive and diverse datasets. By assigning greater weights to the dimensions that predominantly capture negation-related information, they are expected to recognize and process negations more effectively. To ensure the universality across different tasks: this negation-aware weight will only be applied for negation related tasks, while the original unweighted embedding is used for other tasks.
@@
where $a$ is a hyperparameter to control the trade off between negation information and general semantic information in text embeddings. Larger $a$ assigns larger weights on to the dimension with positive contribution to negation awareness. For negation related tasks, the dimension weights are then applied through element-wise multiplication to the original embeddings. This is equivalent to apply a weighted cosine similarity on the original embeddings.
5 Experimental results
5.1 Experiments on simple negation
5.1.1 Datasets
@@ 5.1.2 Experiment Setups
5.1.3 Experimental results
Table 1: Performances of text embeddings on STSB dataset: corr is the correlation between original embedding similarity and the human annotations, $corr\_W_{stsb}$ is the correlation between weighted embedding similarity and the human annotations; acc is the accuracy of original embedding; $acc\_W_{stsb}$ is the accuracy of weighted embedding.
@@ 5.1.3 Experimental results
The experimental results of state-of-the-art text embeddings on STSB dataset are shown in Table 1: acc is the test accuracy of original embedding; $acc\_W_{stsb}$ is the accuracy of weighted embedding where the weight is calculated following Equation 5. Generally speaking, universal text embeddings have better performance than the two baseline contextual text embedding models. The proposed dimension weighting method improves the accuracy across all text embeddings with an average improvement of 4.51%. bge-large-en-v1.5 benefits the most from the proposed method with an improvement of 6.91% while gte-base-en-v1.5 has the smallest improvement of 0.91%. To study if improved negation awareness would hurt the topic information understanding, the correlation between human annotated similarity values (mainly based on topic information) and embedding semantic similarity (before and after applying weights) is also calculated as shown in Table 1: corr is the correlation between original embedding similarity and the human annotations, $corr\_W_{stsb}$ is the correlation between weighted embedding similarity and the human annotations. On average, the semantic similarity based on weighted embeddings have slightly reduced correlation (1% less) with human annotations compared to the original embedding.
@@
5.2 Generalization ability of weights learnt from STSB to SemAntoNeg (more complex negation)
@@
Table 2: The generalization ability of the embedding weights learnt from STSB on SemAntoNeg Benchmark.
@@
Compared to STSB data with simple negation, SemAntoNeg Benchmark is more difficult as it contains negation, antonym and double negation (negation plus antonym). In this section, we evaluate the generalization capability of the simple negation awareness embedding weights, learned from STSB, on the SemAntoNeg Benchmark. The results are presented in Table 2. The overall performance of universal text embeddings (with the mean performance of 58.64%) are much better than that of contextual text embedding baselines (with the mean performance of 33.32%). The random guess performance on SemAntoNeg Benchmark is 33.33%, which indicates that contextual text embeddings is prone to predicting the sentence with the highest lexical overlap while ignoring the negation [43]. On average (across all text embedding models), the embedding weights learnt from STSB improves the performance of the original embedding by 2.41%. This suggests that the weights learnt from STSB are able to generalize to SemAntoNeg Benchmark. However, the performance gains for e5-large-v2 and gte-base-en-v1.5 are less noticeable. In terms of model parameter size, it is observed that larger universal text embeddings demonstrate superior comprehension of complex negation compared to their smaller counterparts.
5.3 Generalization ability of proposed solution on SemAntoNeg Benchmark
Table 3: The generalization ability of proposed solution on SemAntoNeg Benchmark: $acc\_W_{anto}^{K}$ means the performance of embedding weights learnt from K training samples (K = 200, 500, 100). The mean and standard deviations of accuracy evaluated over 10 runs are shown in the table.
@@ 5.3 Generalization ability of proposed solution on SemAntoNeg Benchmark
In this section, the objective is to study whether the proposed embedding re-weighting solution is able to learn from complex negation (negation + antonym) from SemAntoNeg Benchmark. Unlike STSB, SemAntoNeg Benchmark has no predefined train test split. Hence, a stratified repeated random sampling approach is used to achieve a robust estimate of the performance. The random splitting is repeated 10 times, with 32% as training data (1000x4 samples) and 68% as test data (2152x4 samples). To test the data efficiency of the proposed solution, we test 3 different training data sizes: 200, 500 and 1000. The corresponding test performances (mean and standard deviations of accuracy evaluated over 10 runs) are shown in Table 3 as $acc\_W_{anto}^{200}$, $acc\_W_{anto}^{500}$, and $acc\_W_{anto}^{1000}$ respectively.
Generally speaking, the proposed weighting method is able to improve the performance on complex negation tasks across all text embedding models with small size training data. When the training data size is only 200 ($acc\_W_{anto}^{200}$ in Table 3), the universal text embeddings’ performance are improved by 14.52% on average, while the contextual text embeddings’ performance are improved by 11.29%. When increasing the training data size from 200 to 500, the average universal text embeddings’ performance are further improved slightly by 1.34% with reduced standard deviations, while the average contextual text embeddings’ performance are further improved by 0.97%. Wilcoxon-Holm post hoc test [12, 8] with Critical Differences (CD) is also done to have an overall statistical comparison as shown in Figure 3: the proposed solution with three training sizes are all significantly better than the original embeddings, but there is no significant difference between the performance on 500 training datasize and 1000 training datasize (when alpha is 0.00001). Similar to the observations in previous sections: 1. larger universal text embeddings demonstrate superior comprehension of negation compared to their smaller counterparts, 2. universal embeddings benefit more from the proposed solution than contextual embeddings.
@@
5.4 Generalization to Large Language Model based universal text embeddings
@@
The objective in this section is to study whether LLM-based high dimensional task-specific embeddings can be further improved by the proposed dimension weighting method. Three state-of-the-art LLM-based text embeddings are selected: SFR-Embedding-2_R with Mistral-7B [23] as backbone, gte-Qwen2-7B-instruct with Qwen2-7B [49] as backbone and gte-Qwen2-1.5B-instruct with Qwen2-1.5B [49] as backbone. The experimental results are shown in Table 4: Similar to the observations from previous sections, the proposed data efficient solution can improve significantly (around 12% improvement on average) the negation awareness of LLM-based text embeddings with small training data size.
@@
Table 4: The performance of proposed solution with LLM-based embedding models on SemAntoNeg Benchmark: $acc\_W_{anto}^{K}$ means the performance of embedding weights learnt from K training samples (K = 200, 500, 100). The mean and standard deviations of accuracy evaluated over 10 runs are shown in the table.
@@ 5.5 Summary
In this section, STSB dataset with simple negation is used firstly to test the effectiveness of the proposed dimension re-weighting method. The proposed solution is able to improve the accuracy of negation awareness task of both BERT-based universal text embeddings (4.68% improvement) and contextual text embeddings (3.83% improvement). The weights learnt on STSB are subsequently applied to SemAntoNeg Benchmark with complex negation (negation + antonym). The experimental results show that the STSB learnt weights can also improve the complex negation awareness with over 2% improvement on BERT-based universal text embedding models. Furthermore, the proposed solution is used to learn weights from small training data on SemAntoNeg Benchmark (200, 500 and 1000): performances based on all three training datasizes are significantly better than the original embedding (with over 10% improvement) across all BERT-based embedding models, which shows that the proposed solution is robust and data efficient. Finally, we assess the method’s effectiveness on three LLM-based embeddings and show that the proposed method can also improve the negation-awareness of high dimensional task-specific LLM-based embeddings.
Sustainability: The environmental impact of AI solutions has attracted a lot of attention in recent years [7, 9]. Compared to existing solutions, the proposed solution is more lightweight and sustainable. Specifically, our solution requires a small training dataset size (200 to 500 instances), necessitates minimal training time (typically under 5 minutes with a single CPU), and incurs significantly lower computational costs, as it operates effectively without the need for GPU resources. Additionally, since our solution does not modify the model parameters, it eliminates the need for storing fine-tuned versions of the model, thereby further reducing resource consumption.
6 Conclusion
@@
Recent literature has witnessed the fast progress and development of universal text embeddings. However, due to the bias in popular evaluation benchmarks, the negation awareness capacity of these models remains unclear. In this work, we start with a holistic analysis that exposes a critical “negation blindness” in state-of-the-art universal embeddings. To efficiently bridge this universality gap, the negation adapter, a lightweight, data-efficient re-weighting mechanism that restores negation sensitivity is proposed without modifying the model parameters. The proposed solution is able to improve text embedding models’ negation awareness significantly on both STSB with simple negations and SemAntoNeg with complex negations. Furthermore, the proposed solution can also improve the negation awareness of LLM-based task-specific high dimensional embeddings. An additional benefit of the proposed solution is its flexibility: post-hoc re-weighting is applied exclusively to tasks that need strong negation awareness, while other tasks like topic clustering/classification use the original unweighted embedding, which ensures the universality of universal text embeddings across different tasks. Future research aims to enhance the proposed semantic adapter for broader applications, integrating the data-efficient solution with automatic training data generation by LLMs to improve the generalization capabilities of universal text embedding models on novel tasks.